A 0.400 kg bird is located at the point r=3t+4j m from the origin and is moving with a velocity v=5i-2j m/s. What is the angular momentum of this particle about the origin?

Grade:12th pass

1 Answers

Karan Gupta
6 years ago
r (vector) = 3i + 4j, v (vector) = 5i – 2j
 
Angular momentum is mass times the cross product of r and v = m*(3i + 4j) X (5i – 2j)
                                                                                                    = 0.4*(-6k – 20k) = 10.4 (-k)